The 2.1-kg collar travels with negligible friction on the vertical rod under the action of the constant force P = 60 N. If the collar starts from rest at A, determine its speed as it passes point B. The value of R is 2.9 m.Chapter D3, Problem D3/087 x Incorrect The 2.1-kg collar travels with negligible friction on the vertical rod under the actio
Answer #1

R12 a Tension in the string م =be f ازيا Tano = Tan O= } O= Tan (2) FBD Q = 2655655 R - Pcoso of a psing psino Ymg esino doe2) Ploso-mg = ma Golcos (26-50) – (2.1) (9.8)) = 2.1(a). 53.66 - 20160) = (2.1) (a) 33.059 = (201) a =) 15:026=a to find its